# Illicit Promotion on Twitter

### v2

@@ 3.1. The PIP Hunter
To hunt PIPs existing on Twitter, a straightforward method is to inspect every tweet, which however requires unlimited access to the Twitter platform. Instead, our PIP hunter is designed to efficiently search Twitter with keywords that are relevant to PIPs. Therefore, our PIP hunter can not only be used by the Twitter platform for internal inspection, but also serve as an effective tool for any third party to audit illicit promotion on Twitter when there is only a limited access to Twitter data. At a high level, the PIP hunter consists of a cycle of four steps. To start, it searches Twitter with PIP-relevant keywords, which is followed by a binary PIP classifier that takes a multilingual tweet text as the input and decide whether it is a PIP or not. Given PIPs identified, the third step is to evaluate the quality of existing PIP keywords and exclude ones of a low PIP hit rate. Then, the last step is to generate keywords from newly captured PIPs and append them to the keyword set so as to boost the next round of PIP hunting. Below, we present more details step by step.
The tweet crawler. Given PIP-relevant search keywords, a tweet crawler is deployed to query Twitter using its searching APIs  [^2] https://developer.twitter.com/en/docs/twitter-api so as to identify tweets and Twitter accounts that are relevant to PIP keywords. Here, the search keywords are either manually crafted in advance or automatically generated from the last round of hunting. Currently, our tweet crawler supports two types of searching keywords: hashtags and Twitter accounts. And the search strategy varies across the keyword types. For hashtag keywords, the standard Twitter search API will be utilized to retrieve tweets relevant to the given hashtag keyword. Then, when the keyword is a Twitter account, the profile of the account will be retrieved along with its latest tweets up to the crawling time. And the number of tweets to retrieve for each account is configurable and should vary across different deployment strategies. In our case, considering the empirical trade-off between collecting more tweets and avoiding repetitive crawling of the same tweet, we set it as 100, i.e., up to 100 latest tweets will be crawled.
Table 1. The distribution of the ground truth dataset for binary PIP classifiers.
@@
To build up this classifier, several options have been extensively explored. The first option is a combination of feature embedding through TF-IDF and classification through classic algorithms including SVM and Random Forest. The second option also considers only the text elements of a post. However, it is built up through fine-tuning a transformer-based multilingual language model (bert-base-multi-lingual-cased(mul, 2023)) which has achieved state-of-the-art performance in many text classification tasks (Devlin et al., 2018). Besides, many posts especially PIPs have media files attached along with the text elements, and these media files (e.g., images or videos) may have visual elements that are important for deciding whether the respective post is a PIP or not. Therefore, the 3rd option we have explored is a multimodal classifier which takes both the visual modality and the text modality into consideration when classifying a post. Here, the text modality is encoded using aforementioned multilingual language model while the visual input is encoded using a pretrained ResNet-152 model (He et al., 2016).
Table 2. The performance of binary PIP classifiers.
Model Precision Recall F1-Score
Text-Only Transformer 97.25% 96.04% 96.64%
Multi-Modality Transformer 96.56% 97.43% 97.00%
SVM 95.29% 95.30% 95.29%
Random Forest 95.08% 94.77% 94.81%
Labeling the ground truth. To train and evaluate these classification options, a ground truth dataset is collected through an iterative labeling process. Specifically, The labeling process involves two labelers independently annotating samples and resolving conflicts periodically. Also, an iterative process is followed for labeling, involving 1) searching Twitter with PIP-relevant keywords, which gives crawled tweets; 2) labeling a sampled subset of the crawled tweets to update ground-truth; 3) training a weak PIP classifier using the updated ephemeral ground-truth; 4) applying the weak PIP classifier to predict crawled tweets that are unlabeled, identifying both false positive and false negative predictions; and 5) updating ground-truth accordingly. Besides, when a sample is labeled, it is assigned with not only the binary PIP class, but also one of the PIP categories as listed in Table 3. This iterative process continues until no new PIP categories emerge, each PIP category is well represented in ground-truth and the PIP classifier has achieved a good performance when evaluated on the crawled tweets. Furthermore, inter-rater agreement rate over 90% is achieved across all labeling tasks. Particularly, for 1,000 samples from ground-truth, the agreement rate is 99.8%.
The final ground truth dataset consists of of 8,408 PIPs and 4,773 non-PIP posts that are diverse and representative in terms of categories and natural languages. As shown in Table 1, the entire samples in this groundtruth dataset are composed in 10 natural languages while the positive samples (PIPs) belong to 11 distinct categories of illicit goods and services. One thing to note, when composing non-PIP posts, we consider only PIP candidates that are not true PIPs, rather than using regular tweets. This is based on the observation that such PIP candidates tend to sit closer to the decision boundary than regular tweets and can help train a more robust PIP classifier. Besides, despite having more PIPs than non-PIPs in the ground truth, we don’t observe any negative impact on the binary classifier’s performance, which is well demonstrated by evaluation on wild tweets and crawled tweets. Instead, we observe that fewer PIPs can yield comparable performance. For instance, by removing 4K PIPs for balance, a binary PIP classifier trained on this dataset achieved a precision of 96.6% and a recall of 97.4%. More PIPs, on the other hand, help enhance the ground-truth diversity in categories and languages, and are used to train/evaluate the multi-class PIP classifiers as introduced in §3.2.
@@
Evaluation. Our evaluation on the PIP classifiers is three-fold: 1) 5-fold cross-validation upon ground-truth; 2) Evaluation on crawled tweets; 3) Evaluation on wild tweets that are randomly sampled from Twitter Archiving Project (i.e., wild tweets).
Table 2 lists the results of the five-fold cross validation. We can see that the text-only transformer-based classifier has achieved a performance that is comparable to that of the multi-modality one, and in the meantime better than that of the classic classifier. Besides, the multi-modality model has the extra cost of downloading and preprocessing of the involved media files, while classic algorithms are inferior to transformer models in terms of data efficiency, multilingual support, automatic feature engineering, and generalizability. We thus choose the text-only transformer model as the default binary PIP classifier.
We further evaluated the selected PIP classifier on unlabeled tweets as collected by our tweet crawler using PIP-relevant keywords. Given prediction results, 500 positive predictions were sampled out for manual validation along with 500 negative predictions. And the manual validation reveals a precision of 94.20% and a recall of 100%. Then, to further evaluate the generalizability of our binary PIP classifier, we applied it to wild tweets, namely, multiple daily tweet snapshots from the Twitter Archiving Project  [^3] https://archive.org/details/twitterarchive, which archives the tweet stream on a daily basis at a sampling ratio of 1%. Still, given the prediction results, 500 positive predictions and 500 negative ones were sampled out for manual validation, which reveals a precision of 96.8%, a recall of 99.4% and an accuracy of 98.1%.
@@
Filtering existing PIP keywords. Some existing PIP keywords may not always work well in terms of triggering new PIPs, a filtering step is thus applied to filter out such ineffective keywords. This is achieved through a threshold-based filtering. Specifically, a metric named as $RCP_{kw}$ is defined as the ratio of new PIPs among all the posts retrieved for a given keyword $kw$ in the current hunting round, and if a keyword has $RCP_{kw}$ lower than a configurable threshold, then, it will be added to a blocklist and will not be used in the future rounds. However, if a blocked keyword haven’t been used for 4 or more rounds but gets extracted again by the keyword generator, it will be unblocked and added back to the keyword set. The threshold of $RCP_{kw}$ has been tuned as 1% in our deployment, which allows us to gain a good tradeoff between the PIP coverage and the hunting efficiency.
@@
Searching keyword generator. The newly captured PIPs will be further fed into the keyword generator so as to extract new keywords which in turn will be appended to the keyword set for the next round of PIP hunting. Through inspecting PIP posts and conducting manual searching experiments, we find out that hashtags extracted from known PIPs can serve as good keywords in terms of discovering new PIPs. Also, when a Twitter account has ever posted a PIP, it either has many historical PIPs or will post PIPs in the short future. Also, the same account may post PIPs of diverse categories. Upon these observations, we compose the search keywords from two sources. One is to extract all the hashtags from identified PIPs, and such seeds are called hashtag keywords. The other is to collect Twitter accounts (users) that either have ever posted PIPs or have their account profiles detected as PIPs, and such seeds are named as account keywords. Many PIP keywords turn out to be effective in terms of discovering previously unknown PIPs, and some of which can even help identify PIPs that belong to different categories, different natural languages, or different campaigns. For instance, searching with the Chinese hashtag keyword ”广州线下” (GuangZhou In-Person) has identified 16,334 distinct PIPs which belong to 5 categories, are composed in 10 different natural languages, and have 107 distinct contacts extracted. Similar examples include the hashtag keywords ogkush and WeedFactory and account keywords sn889 and 426buds.
@@
Deployment. To jump start our PIP hunter, PIP-relevant keywords were first extracted from the thousands of PIPs in the aforementioned groundtruth dataset. We then deployed this PIP hunter for the Twitter platform between November 1, 2022 and April 23, 2023. During the deployment, each round started with searching Twitter with keywords, and ended with new PIPs and new keywords discovered. Then, the resulting new keywords will be fed into the next round along with existing ones. Also, to avoid non-negligible burden to the Twitter servers, our crawler strictly followed the rate limit policies, would suspend the crawling when a rate limit was reached, and would not restart until the rate limit was cleared up. As the hunting process moved forward, the keyword explosion still emerged and conflicted with our limited access to Twitter, e.g., by March 9, 2023, we have got 1,280,113 distinct keywords (405,932 hashtags and 874,175 accounts). Therefore, a random sampling strategy was then applied to limit the workload of our Twitter crawler to around 60K PIP keywords.
In total, we have scanned over 53 millions of tweets, and discovered 12,401,082 PIPs and 580,530 PIP accounts. Besides, despite being constrained by the limited access to other OSN, we have also verified the existence and the concerning prevalence of PIPs on other three OSNs, as detailed in §4.
@@
The applicability to other OSN platforms. We also evaluate whether this PIP hunter is applicable to other OSNs including Facebook, Youtube, and TikTok. However, only manual evaluation was conducted, due to either strict rate limit or the unavailability of programmable searching APIs. As the result, we observe that PIP-relevant hashtags collected from one OSN (e.g., Twitter) turn out to be applicable to other OSNs, which suggests our hashtag-based searching strategy can likely work for all these OSNs. Besides, when predicting posts collected from OSNs other than Twitter, the binary PIP classifier has achieved a performance that is comparable to that for tweets. We thus believe our PIP hunter is capable to fulfill the task of cross-OSN PIP hunting, and we leave it as a future work to further explore cross-OSN illicit promotion.
@@ 3.2. The PIP Analyzer
The multiclass PIP classifier. Given identified PIPs, a multiclass PIP classifier is designed to group PIPs into one of ten categories of illicit goods and services. And these categories are learned from aforementioned labeling process and are defined according to previous works(Yang et al., 2017b; Du et al., 2016; Yang et al., 2021; Hong et al., 2022; Gomez et al., 2022; Liao et al., 2016b) and Twitter’s policies(twi, 2023c). The full list of these categories is shown in Table 3 along with a short description. For each of these categories, we have observed a reasonable volume of samples when labeling PIPs. We also define another category as others to denote PIPs that don’t fit in well for the aforementioned categories. Then, regarding the naming of these categories, we try our best to make them self-explained while keeping aligned with relevant terms used in previous works (Yang et al., 2017a; Yuan et al., 2018; Cao et al., 2012; Lin, [n. d.]; Zhou et al., 2018; West and Bhattacharya, [n. d.]). Particularly, 8 of the 10 categories are considered as illegal in Twitter safety and cybercrime rules (twi, 2023c). Two exceptions are gambling and surrogacy, likely because they vary significantly in legitimacy across jurisdictions. However, we decide to include them as PIP categories considering multiple factors. Particularly, promotion posts of both categories are often correlated with jurisdictions wherein they are illegal. For instance, 68.29% surrogacy posts are in Chinese while surrogacy in China is prohibited. Besides, multiple previous studies on illicit promotion also consider both categories (Yang et al., 2017a; Yuan et al., 2018; Cao et al., 2012; Lin, [n. d.]; Zhou et al., 2018; West and Bhattacharya, [n. d.]).
To build up this classifier,the PIPs in aforementioned ground truth dataset are reused. Similar to the binary PIP classifier, we explored not only text-only but also multimodal classification. For the text-only classification, we fine-tuned bert-base-multilingual-cased(mul, 2023), a multilingual transformer-based language model. And for the multimodal classification, both the text and the visual modality for each PIP have been taken as the input. To build up such a model, ResNet-152 is used to embed the visual input (i.e., an image), and bert-base-multilingual-cased is used to embed the text input. For both models, 80% of the ground truth dataset are used for training, and the remaining 20% are held out for testing. As listed in Table 4, both have achieved very good performance across a set of well-acknowledged metrics. And we choose the text-only model as the default multiclass PIP classifier.
Table 4. The performance of multiclass PIP classifiers.
Model Precision Recall F1-Score
Text-Only Transformer 98.82% 98.80% 98.80%
Multi-Modality Transformer 96.86% 97.73% 97.27%
@@
Among these contacts, website URLs can be easily identified through regular expression matching. However, many URLs deserve further processing for two factors. On the one hand, some URLs were found to redirect visitors to an IM account, e.g., a Telegram URL https://t.me/{account_id}, and we call such URLs as IM URLs. Such IM URLs will be further processed to extract the respective IM type and IM accounts, which is still achieved by defining and applying regular expressions specific to IM platforms. For instance, WhatsApp’s IM URL pattern is https://wa.me/{phone_number} while it is https://line.me/ti/p/{accountID} for LINE. On the other hand, due to the character limit for each tweet, many URLs were found to have been shortened through popular URL shorteners, e.g., https://bit.ly, and https://tinyurl.com. These shortened URLs will be further visited to recover the true URL of the final landing page. One thing to note, a LINE URL could also be shortened by http://lin.ee/XXXXXXX, and it is http://wa.link/XXXXX for Whatsapp. Then, after the true URL is recovered, the IM contacts will be extracted.
The PIP cluster analyzer. Upon PIPs along with respective Twitter accounts and embedded contacts, it is interesting to further uncover the illicit promotion campaigns underpinning PIPs. To achieve this, a cluster analyzer is designed to group PIPs that likely belong to the same underlying campaign. We first abstract PIPs as an undirected graph. In this graph, each PIP account is abstracted as a node of the account type and the size of this node is defined to be proportional to the number of PIPs that this account has posted. Similarly, each PIP contact is also defined as a node but of the contact type. Still the size of a contact node is proportional to the number of PIPs which contain this contact. Then, two nodes will be connected through an edge if they share one or more PIPs. For instance, if a contact node and an account node is connected, there are one or more PIPs that are posted by the account node and contain the entity of the contact node. Similarly, if two contact nodes are directly connected, it means they have been embedded together in one or more PIPs. Given this PIP graph, the flood filling strategy is applied to identify subgraphs isolated to each other. And the resulting subgraphs are considered as separate promotion clusters.

### v3

@@ 3.1. The PIP Hunter
To hunt PIPs existing on Twitter, a straightforward method is to inspect every tweet, which however requires unlimited access to the Twitter platform. Instead, our PIP hunter is designed to efficiently search Twitter with keywords that are relevant to PIPs. Therefore, our PIP hunter can not only be used by the Twitter platform for internal inspection, but also serve as an effective tool for any third party to audit illicit promotion on Twitter when there is only limited access to Twitter data. At a high level, the PIP hunter consists of a cycle of four steps. To start, it searches Twitter with PIP-relevant keywords, which is followed by a binary PIP classifier that takes a multilingual tweet text as the input and decide whether it is a PIP or not. Given the PIPs identified, the third step is to evaluate the quality of existing PIP keywords and exclude ones with a low PIP hit rate. Then, the last step is to generate keywords from newly captured PIPs and append them to the keyword set so as to boost the next round of PIP hunting. Below, we present more details.
The tweet crawler. Given PIP-relevant search keywords, a tweet crawler is deployed to query Twitter using its searching APIs  [^1] https://developer.twitter.com/en/docs/twitter-api so as to identify tweets and Twitter accounts that are relevant to PIP keywords. Here, the search keywords are either manually crafted in advance or automatically generated by the searching keyword generator introduced soon later. Currently, our tweet crawler supports two types of searching keywords: hashtags and Twitter accounts. For hashtag keywords, the standard Twitter search API will be utilized to retrieve tweets relevant to the given hashtag keyword. Then, when the keyword is a Twitter account, the profile of the account will be retrieved along with its latest tweets up to the crawling time. The number of tweets to retrieve for each account is configurable and should vary across different deployment strategies. In our case, considering the empirical trade-off between collecting more tweets and avoiding repetitive crawling of the same tweet, we set it as 100, i.e., up to 100 latest tweets will be crawled.
Table 1. The distribution of the PIP ground truth dataset.
@@
Classification algorithms. Three classification options are explored to build up the binary PIP classifier. The first option is a combination of feature embedding through TF-IDF and classification through classic algorithms including SVM and Random Forest. The second option also considers only the text elements of a post. However, it is built up through fine-tuning a transformer-based multilingual language model (bert-base-multi-lingual-cased(mul, 2023), which has achieved state-of-the-art performance in many text classification tasks (Devlin et al., 2018). And we name it as Text-Only Transformer (see Appendix C for related works on language models). Besides, many posts especially PIPs have media files attached along with the text elements, and these media files (e.g., images or videos) may have visual elements that are important for deciding whether the respective post is a PIP or not. Therefore, the 3rd option we have explored is a multimodal classifier which takes both the visual modality and the text modality into consideration when classifying a post. We thus name it as Multi-Modality Transformer. Here, the text modality is encoded using the aforementioned multilingual language model while the visual input is encoded using a pretrained ResNet-152 model (He et al., 2016).
Labeling the ground truth. The ground truth dataset is collected through an iterative labeling process involving training weak classifiers, identifying false predictions, and updating the ground truth. Also, across labeling tasks, two labelers independently annotate samples and periodically resolve conflicts. And an inter-labeler agreement rate over 90% is consistently achieved. Please refer to Appendix B.1 to learn more details of the labeling process. The final ground truth dataset consists of 8,408 PIPs and 4,773 non-PIPs that are diverse and representative in categories and natural languages. As shown in Table 1, the dataset are composed in 10 natural languages while the positive samples (PIPs) belong to 11 distinct categories of illicit goods and services.
@@
Three-fold evaluation. Our evaluation on the PIP classifiers is three-fold: 5-fold cross-validation upon ground-truth, evaluation on crawled raw tweets, and evaluation on wild tweets that are randomly sampled from Twitter Archiving Project (i.e., wild tweets). After comparing the performance of all classification options, the text-only transformer is selected as the best choice, which consistently achieves a precision of over 94% and a recall of over 96% across all evaluation settings. Please refer to Appendix B.2 to learn the detailed evaluation results.
@@
Filtering existing PIP keywords. Some existing PIP keywords may not always work well in terms of triggering new PIPs, a filtering step is thus applied to filter out such ineffective keywords. This is achieved through a threshold-based filtering. Specifically, a metric named as $RCP_{kw}$ is defined as the ratio of new PIPs among all the posts retrieved for a given keyword $kw$ in the current hunting round, and if a keyword has $RCP_{kw}$ lower than a configurable threshold, then, it will be added to a blocklist and will not be used in the future rounds. However, if a blocked keyword hasn’t been used for 4 or more rounds but gets extracted again by the keyword generator, it will be unblocked and added back to the keyword set. The threshold of $RCP_{kw}$ has been tuned as 1% in our deployment, which allows us to gain a good tradeoff between the PIP coverage and the hunting efficiency.
@@
Searching keyword generator. The newly captured PIPs will be further fed into the keyword generator so as to extract new keywords which in turn will be appended to the keyword set for the next round of PIP hunting. The search keywords are composed from two sources. One is to extract all the hashtags from identified PIPs, and such seeds are called hashtag keywords. The other is to collect Twitter accounts (users) that either have ever posted PIPs or have their account profiles detected as PIPs, and such seeds are named as account keywords. Many PIP keywords turn out to be effective in terms of discovering previously unknown PIPs, and some can even identify PIPs tha belong to different categories, natural languages, or campaigns. For instance, searching with the Chinese hashtag keyword ”广州线下” (GuangZhou In-Person) has identified 16,334 distinct PIPs which are of 5 categories and 10 different natural languages, and have 107 distinct contacts extracted.
Deployment. To jump start our PIP hunter, PIP-relevant keywords were first extracted from the thousands of PIPs in the aforementioned ground truth dataset. As PIPs in the ground truth are diverse in categories and languages, so are the derived keywords. Regarding the keyword selection, we also observe that, to capture diverse PIPs with high coverage, the keyword set doesn’t have to well cover all categories or languages, nor to be uniformly distributed in categories and languages. For example, when starting to compose the ground truth, we only had a keyword set of less than 10 hashtags relevant to drugs and pornography. However, leveraging this small and skewed set, manual execution of the PIP hunting workflow still led to the discovery of PIPs of 10 categories and diverse languages. Multiple factors have contributed to this snowballing effect, e.g., the same hashtag(especially English ones) can be embedded in PIPs of multiple categories and languages.
@@
We then deployed this PIP hunter for the Twitter platform between November 1, 2022 and April 23, 2023. During the deployment, each round started with searching Twitter with keywords, and ended with new PIPs and new keywords discovered. Then, the resulting new keywords will be fed into the next round along with existing ones. Also, to avoid a non-negligible burden to the Twitter servers, our crawler strictly followed the rate limit policies, would suspend the crawling when a rate limit was reached, and would not restart until the rate limit was cleared up. As the hunting process moved forward, the keyword explosion still emerged and conflicted with our limited access to Twitter, e.g., by March 9, 2023, we got 1,280,113 distinct keywords (405,932 hashtags and 874,175 accounts). Therefore, a random sampling strategy was then applied to limit the workload of our Twitter crawler to around 60K PIP keywords. In total, we have scanned over 53 million tweets and discovered 12,401,082 PIPs and 580,530 PIP accounts. Besides, despite being constrained by the limited access to other OSNs, we have also verified the existence and the concerning prevalence of PIPs on the other three OSNs, as detailed in §4.
@@
The applicability to other OSN platforms. We also evaluated the PIP hunter on Facebook, YouTube, and TikTok, though only manual assessments were possible due to strict rate limits or lack of APIs. Despite these limitations, we observe that PIP-relevant hashtags collected from one OSN (e.g., Twitter) turn out to be applicable to other OSNs, which suggests our hashtag-based searching strategy can likely work for all these OSNs. Besides, when predicting posts collected from OSNs other than Twitter, the binary PIP classifier has achieved a performance that is comparable to that for tweets. We thus believe our PIP hunter is capable of fulfilling the task of cross-OSN PIP hunting, and we leave it as a future work to further explore cross-OSN illicit promotion.
@@ 3.2. The PIP Analyzer
The multiclass PIP classifier. Given identified PIPs, a multiclass PIP classifier is designed to group PIPs into one of ten categories of illicit goods and services. And these categories are learned from aforementioned labeling process and are defined according to previous works(Yang et al., 2017b; Du et al., 2016; Yang et al., 2021; Hong et al., 2022; Gomez et al., 2022; Liao et al., 2016b) and Twitter’s policies(twi, 2023c). The full list of categories is shown in Table 2 along with a short description, while more explanations are presented in Appendix D.1 with regards to the naming and illicitness of these categories.
To build up this classifier, the PIPs in the aforementioned ground truth dataset are reused. Similar to the binary PIP classifier, we explored not only the text-only transformer but also the multimodal transformer. Then, in five-fold cross validation, the multimodal model achieved a precision of 96.86% and a recall of 97.73%, while the text-only model outperformed the multimodal one by 2% in precision and 1.5% in recall. We thus selected the text-only transformer as the default multi-class PIP classifier. For more evaluation results, please refer to Appendix D.2.
@@
Among these 6 contact types, websites can be easily extracted through URL extraction, while WhatsApp and LINE accounts are usually embedded in PIPs as shortened URLs and can thus be recovered by dynamic HTTP visits. For the left three contact types, the extraction is abstracted as a task of multi-class named entity recognition, for which we built up a multi-class classifier to recognize the the BIO tags (beginning, inside, and outside) of these contact types. Please refer to Appendix E to learn more details, e.g., the performance of the NER classifier. Applying this contact extractor to all PIPs revealed over 212K distinct contacts, as detailed in §6.1.
The PIP cluster analyzer. Upon PIPs along with respective Twitter accounts and embedded contacts, it is interesting to further uncover the illicit promotion campaigns underpinning PIPs. To achieve this, a cluster analyzer is designed to group PIPs that likely belong to the same underlying campaign. We first abstract PIPs as an undirected graph. In this graph, each PIP account is abstracted as a node of the account type and the size of this node is defined to be proportional to the number of PIPs that this account has posted. Similarly, each PIP contact is also defined as a node but of the contact type. Still the size of a contact node is proportional to the number of PIPs which contain this contact. Then, two nodes will be connected through an edge if they share one or more PIPs. For instance, if a contact node and an account node is connected, there are one or more PIPs that are posted by the account node and contain the entity of the contact node. Similarly, if two contact nodes are directly connected, it means they have been embedded together in one or more PIPs. Given this PIP graph, the flood filling strategy is applied to identify subgraphs isolated to each other. And the resulting subgraphs are considered as separate promotion clusters. The detailed analysis of PIP clusters is presented in §6.2.